Hello, I ran into a conceptual question in a lecture and it said the following was possible: That simple interest and compound interest can be the same over several periods. How is this so?
Answer #1

Answers page-1 * According to Information 1. The interest may be compounded yearly half-yearly or quarterly based on the intePage - 2 Let the principal amount = 100 then after 5 month the interest is, x5 (simple interest) -2.5 nt Total = 100+25= 102.